# Making Interpretable Discoveries from Unstructured Data: A High-Dimensional Multiple Hypothesis Testing Approach

### v1

@@ 3.1 Setup
Let $n]:=\{1,\dots,n\}$, and let $x\mapsto\log x$ be the natural logarithm. The researcher has access to a dataset of size $n$, $\{(W_{i},Z_{i})\}_{i=1}^{n}$, which is sampled i.i.d. from some (super-)population of interest $P$. The $Z_{i}\in\mathcal{Z}$ are unstructured data instances and the $W_{i}\in\mathcal{W}$ are any other observed covariates of interest (e.g., for one empirical example considered in the following section, $W_{i}=D_{i}\in\mathcal{W}=\{0,1\}$, a binary treatment status indicator). The space $\mathcal{Z}$ is typically high-dimensional and semantically poor, e.g., if each $Z_{i}$ was a $244\times 244$ pixel image, $\mathcal{Z}$ might then be the space of all $3\times 244\times 244$ arrays of RGB values, or if $Z_{i}$ was text that was truncated up to some maximum length, $\mathcal{Z}$ might be the space of all binary matrices of a certain dimensionality, which are concatenated one-hot encodings for each word, term, or token in the text with respect to a specific pre-defined vocabulary.
The researcher has access to the function $\texttt{Dict}:\mathcal{Z}\to\{0,1\}^{p}$, which maps an unstructured data observation to a a sparse binary vector. This function is computed by passing an unstructured data point to a DNN equipped with a dictionary learning model (e.g., a LLM equipped with pretrained SAE when $Z_{i}$ are texts), recording the feature activations from the dictionary model, and then aggregating and binarizing these dictionary feature activations to form a single indicator for each feature in the dictionary $j\in[p]$. There is a large possible space of Dict functions the researcher could implement, though an especially straightforward choice for the use case of text data is implementing a Dict function based on the SAE activations at a single layer of the DNN, for which the $j$-th entry of the output is equal to 1 if the $j$-th dictionary feature activated on any token of the input text, and 0 otherwise:
@@
We denote $Y_{i}:=\mathtt{Dict}(Z_{i})\in\{0,1\}^{p}$ as the output of the dictionary transformation of the unstructured data. As such, each $Y_{ij}\in\{0,1\}$ has the natural interpretation that a particular dictionary feature $j$ activated for the unstructured data instance $i$, and, e.g., $\bar{Y}_{\bullet,j}:=n^{-1}\sum_{i=1}^{n}Y_{ij}$ corresponds to an estimate of the probability of the presence of a feature $j$ in the population $P$ of interest. [^3] Dictionary feature activations are typically positively valued scalars, where magnitude is thought to correspond to some notion of intensity of activation; other works in mechanistic interpretability consider using these raw activation values directly, perhaps with max or average pooling across tokens. However, presently, magnitude of activation is not a well-understood or highly interpretable quantitative property of dictionary learning methods, and as such the default Dict function advocated for in the proposed framework does not incorporate this information. That said, researchers interested in extending this framework may find value in alternative Dict functions that leverage activation intensity.
The crux of the proposed framework is that inference on functionals of $Y_{ij}$ is desirable because:
@@
    Each $Y_{ij}$ is monosemantic, and therefore functionals of $Y_{ij}$ for a given $j$ are interpretable, e.g., $E_{P}[Y_{ij}]$ is the probability of monosemantic feature $j$ appearing in an unstructured data observation in population $P$.
@@
    Collectively, the $p$ features in $Y_{ij}$ for any given $i$ are exhaustive of some vast space of concepts of interest to the researcher; intuitively, they are the set of all monosemantic features that a LLM needed to learn to perform well on next token prediction for a massive, internet-scale corpus of text (or other similarly massive unstructured data source). This means the researcher need not pre-specify any particular concepts of interest, reducing researcher degrees of freedom without limiting discovery.
@@
What is needed, then, for principled inference on functionals of $Y_{ij}$ is a high-dimensional selective inference procedure with control over a generalized error rate. To be as well-powered as possible, we want to focus on testing procedures that employ resampling methods for estimating the true covariance across test statistics, such that conservative protection against worst-case dependence is not required. To keep things as general as possible, we also want to allow for only asymptotically valid test statistics and p-values, as is common to much of econometric analysis. In the following sections, theory and corresponding statistical procedures are developed to achieve exactly these aims, and are stated as generally as possible to accommodate settings of interest even beyond analysis of unstructured data.
To pursue inference, consider defining $p\gg n$ “one-sided” hypotheses $\{H_{0,j}\}_{j\in[p]}$, where
$H_{0,j}:\theta_{j}(P):=E_{P}[X_{ij}]\leq 0,\quad X_{ij}=h(W_{i},Y_{ij}),$
for some measurable function $h$. Two choices of $h$ considered in the empirical examples in this paper include $h(W_{i},Y_{ij})=Y_{ij}$ (testing probabilities of dictionary feature activation) and $h(W_{i},Y_{ij})=\frac{D_{i}-\pi}{\pi(1-\pi)}Y_{ij}$ (testing differences in probabilities of activation across groups under known, fixed group assignment), though many others are possible, lending to the flexibility of the framework. As shorthand, we will denote $X:=\{X_{i}\}_{i=1}^{n}$. Consider also defining a set of “two-sided” hypotheses of interest $\{\tilde{H}_{0,j}\}_{j\in[p]}$ as
$\tilde{H}_{0,j}:\theta_{j}(P):=E_{P}[X_{ij}]=0.$
@@
$T_{n,j}:=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}X_{ij},\quad T_{n}:=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}X_{i}.$
For the purposes of developing high-dimensional central limit theory later on, further define
@@
We will denote $\Sigma:=E\left[S_{n}S_{n}^{\mathtt{T}}\right]=n^{-1}\sum_{i=1}^{n}E[(X_{i}-E[X_{i}])(X_{i}-E[X_{i}])^{\mathtt{T}}]$, which simplifies to $\Sigma=E[(X_{i}-E[X_{i}])(X_{i}-E[X_{i}])^{\mathtt{T}}]$ under i.i.d. data. Further, for any subset of indices $K\subseteq[p]$, we will denote $S_{n,K}:=(S_{n,j}:j\in K)^{\mathtt{T}}$.
3.2 High-Dimensional $k$-FWER Exceedance Control
We construct methods that provide high-dimensional $k$-FWER exceedance control asymptotically, or control of the probability of making $k$ or more rejections of true nulls in the large sample limit when $p$ is growing (much) faster than $n$ in the relevant asymptotic thought experiment. Naturally, $k$-FWER control for $k=1$ is control of the FWER, and larger choices of $k$ permit more discoveries.
To proceed, we will adapt strategies from Romano and Wolf [2007] for achieving large sample $k$-FWER control to the high-dimensional setting. The testing procedures of Romano and Wolf [2007] rely on the asymptotic validity of bootstrap approximations of the distribution of the $k$-th largest coordinate (or “$k$-max”) of $S_{n}$. Thus, a bootstrap procedure for the $k$-max statistic that is asymptotically valid in high-dimensions would allow one to run the algorithms of Romano and Wolf [2007] in the high-dimensional setting with their stated control under minimal modifications.
High-dimensional bootstraps must be supported by appropriate high-dimensional central limit theory. Fortunately, the recent work of Ding et al. [2025] provides high-dimensional central limit theory for the $k$-th largest coordinate of a scaled sum of centered independent random vectors so long as $k$ is very small, i.e., $k$ is fixed in the asymptotic thought experiment. We will build on these “small kk” results. Assume $n\geq 3$ and $p\geq 3$. Let $b_{1}>0$ and $b_{2}>0$ be some constants such that $b_{1}\leq b_{2}$, and let $B_{n}>1$ be a sequence of constants, where it is possible that $B_{n}$ diverges. Then we make the following assumptions.
@@ 3.2 High-Dimensional $k$-FWER Exceedance Control
These are mild conditions on the tails and moments of the data $X_{ij}$, which are stated in a way that accommodates data that are independent but not identically distributed (i.n.i.d.). Part (i) of Assumption M simply requires that $X_{ij}$ be sub-exponential, or, equivalently, have an Orlicz norm in $\Psi_{1}$ bounded by $B_{n}<\infty$. The $B_{n}$ is indexed by $n$ to accommodate growing tail thickness within the sub-exponential regime as $p=p_{n}$ grows in the asymptotic thought experiment with i.n.i.d. data. Part (ii) of Assumption M insists that the second moments of the data be bounded away from zero, appropriately stated for i.n.i.d. data. Part (iii) insists on bounded fourth moments in a similar fashion. As such, for i.i.d. data, the case considered in this paper, we may simply require for some fixed $B<\infty$ and fixed $b_{1},b_{2}$ with $b_{1}\leq b_{2}$: (i) $E\left[\exp\left(\left|X_{ij}\right|/B\right)\right]\leq 2$; (ii) $b_{1}^{2}\leq E[X_{ij}^{2}]$; and (iii) $E[X_{ij}^{4}]\leq B^{2}b_{2}^{2}$. Note that, in many use cases of interest for dictionary learned features, the $X_{ij}$ are both bounded and studentized, aiding the plausibility of these assumptions.
In practice, some fraction of dictionary features may be “dead” for any given dataset, meaning they are degenerate at zero [Bricken et al., 2023]. Of course, these features are not of interest for the purposes of discovery. This phenomenon is tolerated by the framework so long as the count of degenerate features is smaller than any relevant pkp-k, in which case, in any finite sample, the $k$-th largest coordinate of all relevant statistics is invariant to whether or not the degenerate coordinates are dropped. As such, conceptually, one should think of $j$ as ranging over the $p$ non-degenerate features. In practice, to perform studentization, it is recommended that these degenerate coordinates are explicitly filtered out; studentization is discussed further in Section 3.3.
@@
This is the key rate condition needed for (sup-norm) Gaussian or bootstrap approximation error to go to zero asymptotically under the high-dimensional CLTs discussed in both Ding et al. [2025] and Chernozhuokov et al. [2022]. Rewritten, it says that
@@
where $\xi_{i}\overset{iid}{\sim}N(0,1)$ and $\bar{X}_{n}:=n^{-1}\sum_{i=1}^{n}X_{i}$. We also introduce $x\mapsto x_{[k]}$ as notation for the function that selects the $k$-th largest coordinate of a vector $x$, i.e., $x_{[k]}=k\text{-}\!\max(x)$ in the notation of Romano and Wolf [2007]. For the purposes of bootstrapping, we will denote the data-conditional probability measure $P^{B}(\cdot):=P(\cdot\mid X)$.
With these assumptions and notations in place, we now introduce statistical procedures for controlling the $k$-FWER for small $k$, based on the step-wise algorithms of Romano and Wolf [2007].
@@
[High-dimensional $k$-FWER exceedance control for small $k$, one-sided]theoremSmallk Consider the method of Algorithm 2.1 or 2.2 in Romano and Wolf [2007] with test statistics $T_{n}$ of hypotheses $\{H_{0,j}\}_{j\in[p]}$ and critical values $\hat{c}_{n,K}(1-\alpha,k)$ given by the $1-\alpha$ quantile of $S_{n,K,[k]}^{B}$ under $P^{B}$. Assume that $k$ is fixed (i.e., not growing with $n,p$). Then under Assumptions M and R:
@@
    If $H_{0,j}$ is false and $\theta_{j}(P)\gg B_{n}\sqrt{\log p/n}$, then the probability that the step-down method rejects $H_{0,j}$ tends to 1.
@@
[High-dimensional $k$-FWER exceedance control for small $k$, two-sided]theoremSmallktwo Consider the method of Algorithm 2.1 or 2.2 in Romano and Wolf [2007] with test statistics $|T_{n}|$ of hypotheses $\{\tilde{H}_{0,j}\}_{j\in[p]}$ and critical values $\hat{c}_{n,K}(1-\alpha,k)$ given by the $1-\alpha$ quantile of $|S_{n,K}^{B}|_{[k]}$ under $P^{B}$. Assume that $k$ is fixed (i.e., not growing with $n,p$). Then under Assumptions M and R:
@@
    If $\tilde{H}_{0,j}$ is false and $|\theta_{j}(P)|\gg B_{n}\sqrt{\log p/n}$, then the probability that the step-down method rejects $\tilde{H}_{0,j}$ tends to 1.
Importantly, the procedures of Theorems R and R are only valid for small $k$. To see this explicitly, note that the sup-norm bootstrap or Gaussian approximation error in the high-dimensional CLTs for the $k$-th largest coordinate introduced in Ding et al. [2025] only goes to zero if
@@
meaning that $k$ must be fixed or must grow incredibly slowly: for all practical purposes, $k$ needs to be quite small. As such, the theory of Ding et al. [2025] is unsuitable for making progress on FDP exceedance control, which requires a $k$ that may grow linearly with $p$. An important goal of future research would be to establish valid high-dimensional FDP exceedance control using appropriate high-dimensional central limit theory.

### v2

@@ 3.1 Setup
By way of notation, define $n]:=\{1,\dots,n\}$, and let $x\mapsto\log x$ be the natural logarithm. Consider that the researcher has access to a dataset of size $n$, $\{(W_{i},Z_{i})\}_{i=1}^{n}$, which is sampled i.i.d. from some (super-)population of interest represented by distribution $P$. The $Z_{i}\in\mathcal{Z}$ are unstructured data instances and the $W_{i}\in\mathcal{W}$ are any other observed covariates of interest. The space $\mathcal{Z}$ is typically high-dimensional and semantically poor, e.g., if each $Z_{i}$ was a $244\times 244$ pixel image, $\mathcal{Z}$ might then be the space of all $3\times 244\times 244$ arrays of RGB values, or if $Z_{i}$ was text that was truncated up to some maximum length, $\mathcal{Z}$ might be the space of all binary matrices of a certain dimensionality, which are concatenated one-hot encodings for each word, term, or token in the text with respect to a specific pre-defined vocabulary.
Example 1 (Continued).
The $Z_{i}$ is a text response elicited for experimental participant $i$; the $W_{i}\in\mathcal{W}=\{0,1\}$ is a treatment status indicator for experimental participant $i$.
The researcher has access to a function which creates feature dictionaries from unstructured data points, $\texttt{Dict}:\mathcal{Z}\to\{0,1\}^{p}$. This function is computed by passing an unstructured data point to a DNN equipped with a dictionary learning model (e.g., a LLM equipped with pretrained SAE when $Z_{i}$ are texts); recording the feature activations from the dictionary model; and then pooling and binarizing these dictionary feature activations to form a single indicator for each feature in the dictionary, indexed $j\in[p]$. There is a large possible space of Dict functions the researcher could implement, though an especially straightforward choice for the use case of text data is implementing a Dict function based on the SAE activations at a single layer of the DNN, for which the $j$-th entry of the output is equal to 1 if the $j$-th dictionary feature activated on any token of the input text, and 0 otherwise:
@@
We denote $Y_{i}:=\mathtt{Dict}(Z_{i})\in\{0,1\}^{p}$ as the output of the dictionary transformation of the unstructured data: the “feature dictionary.” As such, each “feature indicator” or “dictionary entry” $Y_{ij}\in\{0,1\}$ has the natural interpretation that a particular dictionary feature $j$ activated for the unstructured data instance ii. [^7] Dictionary feature activations are typically positively valued scalars, where magnitude is thought to correspond to some notion of intensity of activation; other works in mechanistic interpretability consider using these raw activation values directly, perhaps with max or average pooling across tokens. However, presently, magnitude of activation is not a well-understood or highly interpretable quantitative property of dictionary learning methods, and as such the default Dict function advocated for in the proposed framework does not incorporate this information. That said, leveraging feature activation intensity is empirically shown to be useful for deriving good autointerpretation generation strategies; see Section 3.4 for more on this.
Example 1 (Continued).
Each text response $Z_{i}$ is passed through an open-source LLM equipped with pretrained SAE. The SAE activations for each token of $Z_{i}$ at the $l$-th layer of the LLM are max-pooled and then thresholded at zero, yielding a $p\approx 10^{5}$-dimensional binary vector $Y_{i}$ for each experimental participant $i$. Each $Y_{ij}\in\{0,1\}$ is a binary indicator of a particular concept being expressed in experimental participant $i$’s text response, e.g., $Y_{i1}$ could indicate whether or not experimental participant $i$’s text response contains phrases indicating uncertainty.
The crux of the proposed framework is that inference on functionals of (the distribution of) each $Y_{ij}$ is desirable because:
@@
    Each $Y_{ij}$ is monosemantic (indicates the presence of a single, intelligible concept) and therefore functionals for $Y_{ij}$ for any given $j$ (and associated hypotheses) are interpretable, e.g., $E_{P}[Y_{ij}]$ is the probability of human-interpretable feature $j$ appearing in an unstructured data observation in population $P$.
@@
    Collectively, the $p$ features in $Y_{i}$ for any given $i$ are exhaustive of some vast space of concepts of interest to the researcher. Heuristically, they encode the set of all concepts that a LLM needed to learn to perform well on next token prediction for a massive, internet-scale corpus of text (or other similarly massive unstructured data source). This means the researcher need not specify ex-ante or ex-post any particular concepts of interest, reducing the conceptual and statistical complications associated with data snooping without limiting discovery.
@@
What is needed, then, for principled inference on functionals of $Y_{ij}$ is a high-dimensional selective inference procedure with control over a generalized selective error rate. To be as well-powered as possible, we want to focus on testing procedures that employ resampling methods for estimating the true covariance across test statistics, such that conservative protection against worst-case dependence is not required. To keep things as general as possible, we also want to allow for only asymptotically valid test statistics and p-values, as is common in much of econometric analysis. In the following sections, theory and corresponding statistical procedures are developed to achieve exactly these aims, and are stated as generally as possible to accommodate settings of interest even beyond analysis of unstructured data.
To set up multiple hypothesis testing, consider defining $p\gg n$ “one-sided” feature/concept-level hypotheses $\{\tilde{H}_{0,j}\}_{j\in[p]}$, where
$\tilde{H}_{0,j}:\theta_{j}(P):=E_{P}[X_{ij}]\leq 0,\quad X_{ij}=h(W_{i},Y_{ij}),$
for some measurable function $h$. Consider also defining a set of “two-sided” hypotheses of interest $\{H_{0,j}\}_{j\in[p]}$ as
$H_{0,j}:\theta_{j}(P):=E_{P}[X_{ij}]=0.$
As shorthand, we will denote $X:=\{X_{i}\}_{i=1}^{n}$. Many useful choices of $h$ are possible under this framework, lending to its flexibility.
Example 1 (Continued).
To infer average treatment effects on dictionary features, for treatment indicator $W_{i}$, fixed probability of treatment $\pi$ (bounded away from zero and one), and feature indicator potential outcomes $Y_{ij}(w)$ under SUTVA,
$H_{0,j}:E[Y_{ij}(1)-Y_{ij}(0)]=E[X_{ij}]=E\left[\frac{W_{i}-\pi}{\pi(1-\pi)}Y_{ij}\right]=0.$
In other words, we set $X_{ij}=h(W_{i},Y_{ij}):=\frac{W_{i}-\pi}{\pi(1-\pi)}Y_{ij}$, a Horvitz–Thompson transformation for computing a difference in means.
@@
$T_{n,j}:=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}X_{ij},\quad T_{n}:=\frac{1}{\sqrt{n}}\sum_{i=1}^{n}X_{i},$
(and taking absolute values, as appropriate). For the purposes of developing high-dimensional central limit theory later on, further define
@@
We will denote $\Sigma:=E\left[S_{n}S_{n}^{\mathtt{T}}\right]=n^{-1}\sum_{i=1}^{n}E[(X_{i}-E[X_{i}])(X_{i}-E[X_{i}])^{\mathtt{T}}]$, which simplifies to $\Sigma=E[(X_{i}-E[X_{i}])(X_{i}-E[X_{i}])^{\mathtt{T}}]$ under i.i.d. data. Further, for any subset of indices $K\subseteq[p]$, we will denote the subvector $S_{n,K}:=(S_{n,j}:j\in K)^{\mathtt{T}}$.
3.2 High-Dimensional $k$-FWER Control
We construct methods that provide high-dimensional $k$-FWER control asymptotically, i.e., control of the probability of making $k$ or more rejections of true nulls in the large sample limit when $p$ is growing (much) faster than $n$ in the relevant asymptotic thought experiment. Naturally, $k$-FWER control for $k=1$ is control of the FWER, and larger choices of $k$ make the permit more discoveries.
To proceed, we will adapt strategies from Romano and Wolf [2007] for achieving large sample $k$-FWER control to the high-dimensional setting. The testing procedures of Romano and Wolf [2007] rely on the asymptotic validity of bootstrap approximations of the distribution of the $k$-th largest coordinate (or “$k$-max”) of $S_{n}$. Thus, the core intuition for results stated in this section is that, by building on a bootstrap procedure for the $k$-max statistic that is asymptotically valid in high-dimensions, we may salvage the general construction of the algorithms of Romano and Wolf [2007] to achieve analogous generalized selective error guarantees in high-dimensions to the low-dimensional setting originally considered by them.
High-dimensional bootstraps must be validated by appropriate high-dimensional central limit theory. Fortunately, the recent work of Ding et al. [2025] provides high-dimensional central limit theory for the $k$-th largest coordinate of a scaled sum of centered independent random vectors so long as $k$ is very small, i.e., $k$ is fixed in the asymptotic thought experiment. We will build on these “small kk” results. Assume $n\geq 3$ and $p\geq 3$. Let $b_{1}>0$ and $b_{2}>0$ be some constants such that $b_{1}\leq b_{2}$, and let $B_{n}>1$ be a sequence of constants, where it is possible that $B_{n}$ diverges. Then we make the following assumptions.
@@ 3.2 High-Dimensional $k$-FWER Control
These are mild conditions on the tails and moments of the data $X_{ij}$, which are stated in a way that accommodates data that are independent but not identically distributed (i.n.i.d.). Part (i) of Assumption M simply requires that $X_{ij}$ be sub-exponential, or, equivalently, have an Orlicz norm in $\Psi_{1}$ bounded by $B_{n}<\infty$. The $B_{n}$ is indexed by $n$ to accommodate growing tail thickness within the sub-exponential regime as $p=p_{n}$ grows in the asymptotic thought experiment with i.n.i.d. data. Part (ii) of Assumption M insists that the second moments of the data be bounded away from zero, appropriately stated for i.n.i.d. data. Part (iii) insists on bounded fourth moments in a similar fashion. As such, for i.i.d. data, the case considered in the applications of this framework, we may simply require for some fixed $B<\infty$ and fixed $b_{1},b_{2}$ with $b_{1}\leq b_{2}$: (i) $E\left[\exp\left(\left|X_{ij}\right|/B\right)\right]\leq 2$; (ii) $b_{1}^{2}\leq E[X_{ij}^{2}]$; and (iii) $E[X_{ij}^{4}]\leq B^{2}b_{2}^{2}$. Note that, in many use cases of interest for dictionary learned features, the $X_{ij}$ will be both bounded (almost surely) and studentized, making these assumptions highly plausible.
[…]
@@
This is the key rate condition needed for (sup-norm) Gaussian or bootstrap approximation error to go to zero asymptotically under the high-dimensional CLTs discussed in both the works of Ding et al. [2025] and Chernozhuokov et al. [2022]. Rewritten, it says that
@@
Both Assumptions M and R are the state of the art in the high-dimensional central theory from which this framework draws [Chernozhuokov et al., 2022, Ding et al., 2025]. Indeed, these conditions are the exact same as those used in both Chernozhuokov et al. [2022] and Ding et al. [2025]. Notably, these assumptions do not require non-degeneracy of $\Sigma$, a well-established and beneficial feature of these results [Chernozhukov et al., 2023]. Under assumptions of non-degenerate $\Sigma$, improved rate conditions may be used in place of Assumption R.
@@
where $\xi_{i}\overset{iid}{\sim}N(0,1)$ and $\bar{X}_{n}:=n^{-1}\sum_{i=1}^{n}X_{i}$. We also introduce $x\mapsto x_{[k]}$ as notation for the function that selects the $k$-th largest coordinate of a vector $x$, and use notation $x\mapsto k\text{-}\!\max(x)$ to denote the $k$-the largest element of a set, per the notation of Romano and Wolf [2007]. For the purposes of bootstrapping, we will denote the data-conditional probability measure $P^{B}(\cdot):=P(\cdot\mid X)$.
With these assumptions and notations in place, we now introduce statistical procedures for controlling the $k$-FWER for small $k$ in large samples, based on the step-wise algorithms of Romano and Wolf [2007].
@@
[High-dimensional $k$-FWER control for small $k$, one-sided]theoremSmallk Consider the method of Algorithm 2.1 or 2.2 in Romano and Wolf [2007] with test statistics $T_{n}$ of hypotheses $\{\tilde{H}_{0,j}\}_{j\in[p]}$ and critical values $\hat{c}_{n,K}(1-\alpha,k)$ given by the $1-\alpha$ quantile of $S_{n,K,[k]}^{B}$ under $P^{B}$. Assume that $k$ is fixed (i.e., not growing with $n,p$). Then under Assumptions M and R:
@@
    If $\tilde{H}_{0,j}$ is false and $\theta_{j}(P)\gg B_{n}\sqrt{\log p/n}$, then the probability that the step-down method rejects $\tilde{H}_{0,j}$ tends to 1.
This result is the natural high-dimensional analog of Theorem 3.1 (as well as Theorem 3.3) of Romano and Wolf [2007]. It says that, so long as $k$ is small (in the relevant asymptotic sense), Algorithm 2.1 or Algorithm 2.2 of Romano and Wolf [2007] may be used to asymptotically achieve control of the $k$-FWER at any desired level $\alpha$ in high-dimensional settings, and moreover that these algorithms are consistent against local alternatives that are not shrinking towards zero too quickly [^10] Under Assumption R, note that $B_{n}\sqrt{\log p/n}=o(1)$. (i.e., intuitively, in sufficiently large samples, we are able to appropriately reject a false null so long as $\theta_{j}(P)=E_{P}[X_{ij}]$ is not very close to zero). The proof of this result is available in Appendix Section 6.4, and, as previously alluded to, relies on the insight that the low-dimensional $k$-FWER controlling algorithms of Romano and Wolf [2007] can be adapted to and analyzed in the high-dimensional setting using new results in high-dimensional central limit theory for order statistics (e.g., Ding et al. [2025]). In the following theorem, we state the sibling procedure for high-dimensional control of the $k$-FWER in two-sided testing settings; it is the high-dimensional analog of Theorem 3.2 of Romano and Wolf [2007].
@@
[High-dimensional $k$-FWER control for small $k$, two-sided]theoremSmallktwo Consider the method of Algorithm 2.1 or 2.2 in Romano and Wolf [2007] with test statistics $|T_{n}|$ of hypotheses $\{H_{0,j}\}_{j\in[p]}$ and critical values $\hat{c}_{n,K}(1-\alpha,k)$ given by the $1-\alpha$ quantile of $|S_{n,K}^{B}|_{[k]}$ under $P^{B}$. Assume that $k$ is fixed (i.e., not growing with $n,p$). Then under Assumptions M and R:
@@
    If $H_{0,j}$ is false and $|\theta_{j}(P)|\gg B_{n}\sqrt{\log p/n}$, then the probability that the step-down method rejects $H_{0,j}$ tends to 1.
To comment explicitly on why the procedures of Theorems R and R are only valid for small $k$, note that the sup-norm bootstrap or Gaussian approximation error in the high-dimensional CLTs for the $k$-th largest coordinate introduced in Ding et al. [2025] only goes to zero if
@@
meaning that $k$ must be fixed or must grow incredibly slowly: for all practical purposes, $k$ needs to be quite small. As such, the theory of Ding et al. [2025] is unsuitable for making progress on FDP exceedance control, which requires a $k$ that may grow linearly (“quickly”) with $p$. A useful goal of future research would be to establish valid high-dimensional FDP exceedance control using further extensions of high-dimensional central limit theory. However, as will be observed in the empirical applications of this paper, small $k>1$ are often sufficient to power meaningful discoveries in relevant social science settings at conventional choices of $\alpha$. [^11] For applied researchers looking for guidance on choice of $k$, note that the given approximation error rate suggests a choice of $k$ on the order of $\log(pn)$.